OpenStudy (chillhill):

Find the constant term in the expansion of (2x^2-[1/x])^6

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle\left(ax^2+b/x\right)^n }\) \(\large\color{#000000 }{ \displaystyle\left({~[ax^3+b]\cdot [1/x]~}\right)^n }\) \(\large\color{#000000 }{ \displaystyle \left(1/x\right)^n\left(ax^3+b\right)^n }\) So, \(\large\color{#000000 }{ \displaystyle\left(ax^3+b\right)^n }\) Will have one constant term, since everything is multiplied times \(\large\color{#000000 }{ \displaystyle \left(1/x\right)^n=x^{-n} }\) You need to find the term of \(\large\color{#000000 }{ \displaystyle\left(ax^3+b\right)^n }\) that will have a power of n (so that the product of powers gives a 0 power)

OpenStudy (solomonzelman):

In your case, \(\large\color{#000000 }{ \displaystyle\left(2x^2+1/x\right)^6 }\) \(\large\color{#000000 }{ \displaystyle\left({~[2x^3+1]\cdot [1/x]~}\right)^6 }\) \(\large\color{#000000 }{ \displaystyle \left(1/x\right)^6\left(2x^3+1\right)^6 }\) So, \(\large\color{#000000 }{ \displaystyle\left(2x^3+1\right)^6 }\) Will have one constant term, since everything is multiplied times \(\large\color{#000000 }{ \displaystyle \left(1/x\right)^6=x^{-6} }\) You need to find the term of \(\large\color{#000000 }{ \displaystyle\left(2x^3+1\right)^6 }\) that will have a power of 6 (so that the product of powers gives a 0 power)

OpenStudy (solomonzelman):

My route seems ong, but you just need to find a single term from (2x^3+1)^6 that will have a power (of x) of 6, and then everything is multiplied times x^(-6), so multiply this term times x^(-6) and you get your constant term.

OpenStudy (solomonzelman):

You can use pascal's triange.

OpenStudy (solomonzelman):

In terms of a and b, I find that this term with power of 6, is: \(\large\color{#000000 }{ \displaystyle 15a^2b^4x^6 }\)